# 3D Mesh Editing using Masked LRMs

### v1

@@ 4.1 Quantitative Comparisons
Table 1 shows novel-view synthesis metrics of our method when compared to MeshLRM. Since our main goal is to edit existing shapes and not to completely generate shapes from scratch, we choose to train our model by randomly selecting 6-8 input views along with one conditional view, giving our LRM a denser input than MeshLRM. We show metrics for both 6 and 8 input views and we compute those on another set of 10 camera poses, different from the input poses. Our method is competitive with the state-of-the-art model on reconstruction, achieving a 2.56 PSNR improvement on the ABO dataset, and a comparable PSNR on GSO. We observe the same phenomenon in perceptual quality measured by LPIPS, where our method significantly outperforms on ABO shapes, and is comparable on GSO shapes. As expected, using 6 views under-performs using 8 views, but only by a slight margin. Our model achieves performance on par with SoTA on reconstructing a diverse set of output poses, indicating that it has learned to effectively “in-paint" the randomly occluded regions in the input views using context from the available unoccluded signal. Since our end goal is mesh shape editing, it is not critical that we surpass the reconstruction quality of prior works, as we only need to ensure a high quality baseline for the output geometry. We further demonstrate qualitatively in Sec. 4.3 that our model indeed learns to inpaint using the conditional signal, instead of only the context from multi-view images, thereby accomplishing feed-forward shape editing through a single view.
4.2 Qualitative Evaluations
Using a bird mesh generated from a text-to-3D model, and several editing targets, we compare our method against other 3D editing methods. Full results are shown in Figure 4. We define a masked region to edit on the head of the bird (omitted from the figure for brevity). Conditional signals provided to our method generated by masked diffusion are shown in the $1^{st}$ row, and our results are shown in the last row.
Table 1: Quantitative Evaluation: We evaluate our model using shapes outside of its training set and compare it to the state-of-the-art LRM. We do so using the ABO and GSO shape datasets, reconstructing the meshes from 6 and 8 posed images. Although our main objective is not direct reconstruction of new shapes, and our masking procedure introduces additional difficulty in the task, we still achieve better than SoTA metrics on the ABO dataset, and comparable to SoTA metrics on the GSO dataset.
@@ 4.2 Qualitative Evaluations
Figure 5: Mesh Editing: The inputs comprise front/back views of the source mesh $1^{st}col$ and a front image used as the conditional view. The $2^{nd}col$ shows the masked area rendered from the front (inset) and a 2D edit. The last column shows our generated mesh from the front/back views (Inset: colors from volumetric rendering).
Optimization methods. In the $2^{nd}$ and $3^{rd}$ rows of Figure 4, we show the results of two text-based mesh optimization methods. Instead of using the edited images themselves, we use the text prompts that we passed to the diffusion model as guidance. The first optimization-based method we compare to ($2^{nd}$ row) is TextDeformer [25] which uses a Jacobian representation [1] for deformations to constrain the edits instead of explicit localization. We observe that TextDeformer struggles with the highly localized nature of our task and globally distorts the mesh, failing to produce an output of acceptable quality in all examples.
The other optimization method we compare to is MagicClay [5], which optimizes both an SDF and a triangle-mesh representation. Note that MagicClay optionally uses a manual “seed" edit so that the optimization task is easier. However, since this requires an additional layer of modeling expertise (i.e. a manual user intervention) our method does not require, we opt out of this step for our comparison. In contrast to TextDeformer, MagicClay relies on selecting a subset of vertices to deform to combat noisy SDS [59] gradients. Since we have a 3D mask, we simply choose the vertices that lie within that region. Although this selection serves to localize the editing process, we observe that the deformations are still noisy. In some cases such as the flower ($4^{th}$ column) and rabbit ears ($5^{th}$ column), the editing process produces semantically correct edits. In other cases such as the fedora ($3^{rd}$ column) and top hat ($6^{th}$ column), the optimization process collapses completely. We note that in both cases, noisy gradients from text-guidance result in results in optimizations that are both unpredictable and uncontrollable. In contrast, the output of our LRM is highly predictable from the selected conditional view, which may be re-generated until desirable or even manually edited.
[…]
@@
Figure 6: Genus changes: Our method can perform genus-changing edits like adding a handle or a hole to the original vase. We show the output of our model from two opposing views in the third column.
@@
InstantMesh. InstantMesh [83] is an LRM-based pipeline that can be used for editing. It relies on using Zero-123++ [68] to generate multi-view images from a single view and then passing these images to an LRM. We compare our results to two versions of InstantMesh. The first version, shown in the $4^{th}$ row of Figure 4, simply passes the single edited view directly to InstantMesh. This results in a poorly reconstructed shape that is particularly thin when compared to the ground truth and the output quality suffers due to the inability of Zero-123++ [68] to generate faithful multi-view images, as discussed in Section 3.4. Methods that rely directly on a separate diffusion module to generate the LRM inputs will run the risk of generating artifacts from inaccurate multiview generation. In comparison, our method does not suffer any such reconstruction artifacts since it uses trivially consistent ground truth renders as the main LRM inputs. The second version of InstantMesh, shown in the $5^{th}$ row, applies the masking procedure from Differential Diffusion [41] to the forward pass of Zero-123++ [68]. Once again, since we define a editing region in 3D, we can restrict the diffusion process to the pixels corresponding to this region. In this version, we observe that the reconstruction is higher quality. However, artifacts still remain, especially near the edited region, due to artifacts generated by Zero-123++.
@@ 4.3 Mesh Editing Characteristics
In Figures 5 and 6, we show more mesh editing examples demonstrating the capabilities of our method. The $1^{st}$ column shows the source mesh rendered from two different viewpoints. The $2^{nd}$ column shows the edited conditional image with a render of the original masked region inset. Our LRM accepts the edited view along with a set of occluded ground-truth renders (omitted from the figure), and predicts an SDF. The last column shows the marching-cubes mesh extracted from the output while the insets depict volumetric renders of the predicted SDF.
Table 2: Runtime Comparison: Our method is significantly faster than optimization methods as it is feed-forward and also faster than InstantMesh which must run Zero-123++.
Optimization-based LRM-based
TextDeformer [25] MagicClay [5] InstantMesh [83] Ours
[…]
Expressiveness. The edits throughout this paper show the expressiveness of our method. The meshes used in Figures 1 and 4 as well as in row 1 of Figure 5 are examples of non-standard shapes – a unique bird mesh generated from a text-to-multiview diffusion model [73] and a “Tele-alien" from the COSEG [79] dataset. Despite being novel shapes, our model is able to give the alien a staff and the bird a hat. The other four rows of 5 consist of edits that are “unnatural" – creating an avocado backrest, replacing the body of a crab with a pineapple, giving a panda wings, and giving a donkey a turtle shell. In every example, our method successfully translates the 2D edit into geometry in a realistic manner. The edits in Figure 6 show a critical benefit of our method. Since the final mesh is constructed entirely from the output of a neural network, there are no geometric restrictions on the type of edit we can do. The last two rows demonstrate the ability of our network to change the genus of a shape by adding a handle or a hole through the middle, which would be impossible for geometry optimization-based methods.
@@ 4.4 Ablations Studies & Discussion
We investigate the effect of using geometric losses during training. Figure 7 compares three LRMs: a model trained by the pipeline described in Sec. 3, a model trained with no geometric supervision, and a model trained by replacing the normal map loss with a depth map loss. We observe that using no geometric supervision results in poor surface quality, highlighted in the red boxes. Since the main training objective is multi-view image reconstruction, the model easily hallucinates correct geometry using colors, without producing an accurate surface. Supervising the predicted depth somewhat mitigates this issue, but the effect is weak and the surfaces are still incomplete. Normal map supervision gives high quality surfaces as shown in the green boxes.
Runtime: In Table 2 we provide performance comparisons between our approach and 3 top-performing recent works all measured on a single A100 GPU. Our method is not only much faster than both optimization-based approaches [5, 25] as it requires only one forward pass, but it also outperforms LRM approaches such as InstantMesh [83] that make use of a multi-view generation model [68].
Limitations & Future Work: Our method is constrained by the expressiveness of editing in the canonical view. While text-to-image models can create a wide range of results, capturing a specific idea may require significant iteration. Furthermore, while our model accurately reconstructs geometry, it replaces the triangulation with one derived from Marching Cubes, which may be undesirable. Finally, our model is limited by the reconstruction quality of modern LRMs which makes performing edits that require extremely intricate details challenging. Future work could focus on improving localization by developing techniques to merge the existing triangulation with the edited output.
@@ 5 Conclusion
In this paper we introduced a new method to perform 3D shape editing. Our work builds upon the recent progress of LRMs by introducing a novel multi-view input masking strategy during training. Our LRM is trained to “inpaint" the masked region using a clean conditional viewpoint to reconstruct the missing information. During inference, a user may pass a single edited image as the conditional input, prompting our model to edit the existing shape in just one forward pass. We believe our method is an significant advancement in shape editing, allowing users to create accurate and controllable edits without 3D modeling expertise.
@@ 6 Introduction
7 Masking Ablation
In order to justify our masking strategy, we train our model masking patches uniformly randomly instead of using 3D occlusions. Figure 8 compares meshes extracted from a model trained using our 3D masking versus masking $25\%$ of patches uniformly at random. We observe that uniformly random patch masking can still generate “roughly correct" shapes, especially adding a moustache to the face in the last row. This is because we add camera pose embeddings after masking, so the model can differentiate masked and non-masked tokens, regardless of their distribution within the image. Furthermore, the reconstruction outside of the masked region is still accurate. However, there still exists a train-test gap between random patches and contiguous patches created by selecting an editing region, which causes significant artifacts in the other three examples. In the first and second rows, we observe a blurring artifact, where the model cannot generate sharp features in the horns on the bird and between the slats of the chair. In the third row, using random patches causes the shape of the turtle shell to be malformed. In comparison, using our masking method produces accurate and sharp geometry in all examples.
8 Comparison to Tailor3D
@@ 9 Comparison to Nerfiller
10 Additional Implementation Details
@@
We trained our models using 64 H100 GPUs with 80GB of RAM each. We use an AdamW optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$ and a weight decay of $0.01$. During stage 1 of training, we train for $30$ epochs. For each batch consisting of $12$ shapes, we randomly sample the number of input views for the batch uniformly at random between $6$ and $8$, not including the $1$ view for the conditional view. We use another $4$ views for supervision. Over the first 1500 iterations, we linearly warm up to a peak learning rate of $44e-$ 4 and then use cosine learning rate decay. During stage 2, to account for increased rendering costs, we reduce the batch size to $8$ shapes. We train for $20$ epochs, with a peak learning rate of $65e-$ 6.
11 Additional Qualitative Examples

### v2

@@ 4.1 Quantitative Comparisons
Table 1 shows novel-view synthesis metrics of our method when compared to InstantMesh and MeshLRM. Since our main goal is to edit existing shapes and not to completely generate shapes from scratch, we choose to train our model by randomly selecting 6-8 input views along with one conditional view, giving our LRM a denser input than MeshLRM. We show metrics for both 6 and 8 input views and we compute those on another set of 10 camera poses, different from the input poses. Our method is competitive with the state-of-the-art model on reconstruction, achieving a 2.56 PSNR improvement on the ABO dataset, and a comparable PSNR on GSO. We observe the same phenomenon in perceptual quality measured by LPIPS, where our method significantly outperforms on ABO shapes, and is comparable on GSO shapes. As expected, using 6 views under-performs using 8 views, but only by a slight margin. Furthermore, our method significantly outperforms InstantMesh. This is to be expected, as InstantMesh infers everything from a single view, while both MeshLRM and MaskedLRM access multi-view information. Our model achieves performance on par with SoTA on reconstructing a diverse set of output poses, indicating that it has learned to effectively “in-paint” the randomly occluded regions in the input views using context from the available unoccluded signal. Since our end goal is mesh shape editing, it is not critical that we surpass the reconstruction quality of prior works, as we only need to ensure a high quality baseline for the output geometry. We further demonstrate qualitatively in Sec. 4.3 that our model indeed learns to inpaint using the conditional signal, instead of only the context from multi-view images, thereby accomplishing feed-forward shape editing through a single view.
Table 1: Quantitative Evaluation: We evaluate our model using test-set shapes and compare it to the state-of-the-art LRM and InstantMesh on the ABO and GSO shape datasets, reconstructing the meshes from 6 and 8 posed images. Despite direct reconstruction of new shapes not being our main goal, and our masking introducing extra difficulty in the task, we still achieve better than SoTA metrics on ABO, and comparable to SoTA metrics on the GSO dataset.
@@
4.2 Qualitative Evaluations
Using a bird mesh generated from a text-to-3D model, and several editing targets, we compare our method against other 3D editing methods. Full results are shown in Figure 5. We define a masked region to edit on the head of the bird (omitted from the figure for brevity). Conditional signals provided to our method generated by masked diffusion are shown in the $1^{st}$ row, and our results are in the last row.
Optimization methods. In the $2^{nd}$ and $3^{rd}$ rows of Figure 5, we show the results of two text-based mesh optimization methods. Instead of using the edited images themselves, we use the text prompts that we passed to the diffusion model as guidance. The first optimization-based method we compare to ($2^{nd}$ row) is TextDeformer [27] which uses a Jacobian representation [1] for deformations to constrain the edits instead of explicit localization.TextDeformer struggles with the highly localized nature of our task and globally distorts the mesh, failing to produce an output of acceptable quality in all examples. We also compare with MagicClay [6], which optimizes both an SDF and a triangle-mesh representation. It also optionally uses a manual “seed” edit so that the optimization task is easier. However, since this requires an additional layer of modeling expertise (i.e. a manual user intervention) our method does not require, we opt out of this step for our comparison. Unlike TextDeformer, MagicClay selects a subset of vertices to deform to combat noisy SDS [64] gradients. Since we have a 3D mask, we simply choose the vertices that lie within that region. Although this selection serves to localize the editing process, we observe that the deformations are still noisy. While sometimes MagicClay edits are semantically correct (flower and rabbit ears), in other cases such as the fedora ($3^{rd}$ column) and top hat ($6^{th}$ column), the optimization process collapses completely. In both cases, noisy gradients from text-guidance result in results in optimizations that are both unpredictable and uncontrollable. In contrast, the output of our LRM is highly predictable from the selected conditional view, which may be re-generated until desirable or even manually edited.
@@ 4.2 Qualitative Evaluations
Figure 4: Genus changes: Our method unlocks genus-changing edits like adding a handle or a hole to the original vase. We show the output of our model from 2 opposing views in the $3^{rd}$ column.
@@
LRM-based Methods. We compare our method against recent top-performing methods that combine multi-view diffusion and reconstruction models. InstantMesh [90] is one such pipeline that may be used for shape editing. It relies on using Zero-123++ [75] to generate multi-view images from a single view and then passing these images to an LRM. To edit, we simply pass an edited image of the original shape As shown in the $4^{th}$ row of Figure 5, this results in a poorly reconstructed shape that is particularly thin when compared to the ground truth and the output quality suffers due to the inability of Zero-123++ [75] to generate faithful multi-view images, as discussed in Section 3.4. Methods that rely directly on a separate diffusion module to generate the LRM inputs will run the risk of generating artifacts from inaccurate multiview generation. In comparison, our method does not suffer any such reconstruction artifacts since it uses trivially consistent ground truth renders as the main LRM inputs.
Figure 5: Mesh Editing Comparisons: Given a mesh (top left) and various image edits as guidance we demonstrate that our approach is the only one that generates multi-view consistent shape edits that follow the guidance. Colors omitted to clearly visualize the edited geometry.
In Figure 6 we also compare to two concurrent works namely PrEditor3D [24] and Instant3Dit [5] that tackle localized editing using multi-view diffusion via text prompting. PrEditor3D [24] performs multi-view editing by first inverting a multi-view diffusion model to obtain Gaussian noise images, and then using the forward process to edit the desired region with a separate text prompt. While PrEditor3D generates semantically correct edits based on the prompts, it produces undesirable artifacts in several examples. In particular, many of the edits lack detail, such as the bunny ears in the top left and the wings in the bottom right. It also fails to produce a pineapple body in the bottom left. Instant3Dit [5] uses masking in multi-view diffusion training instead of LRM training. They use their text-prompted multi-view inpainting model to edit and then use an LRM generate an edited mesh. Similar to PrEditor3D, it produces semantically correct edits that lack realism due to artifacts. Instead of producing sharp bunny ears, Instant3Dit is only capable of adding vague pointed structures to the bird. In the second column, the flower it generates is plant-like but unrealistic. In the bottom row, we see that the pineapple and wings are again semantically correct but lacking in detail.
@@ 4.3 Mesh Editing Characteristics
In Figures 1 and 4, we show mesh editing examples demonstrating the capabilities of our method. The $1^{st}$ column shows the source mesh rendered from different viewpoints. The $2^{nd}$ column shows the edited conditional image with a render of the original masked region inset. Our LRM accepts the edited view along with a set of occluded ground-truth renders (omitted from the figure) and predicts an SDF. The last column shows the mesh extracted from the output while the insets depict volumetric renders of the predicted SDF.
Expressiveness. The edits throughout this paper show the expressiveness of our method. The meshes used in Figures 5 as well as in row 1 of Figure 1 are examples of non-standard shapes – a unique bird mesh generated from a text-to-multiview diffusion model [80] and a “Tele-alien” from the COSEG [86] dataset. Despite being novel shapes, our model is able to give the alien a staff and the bird a hat. The other four rows of 1 consist of edits that are “unnatural” – creating an avocado backrest, replacing the body of a crab with a pineapple, giving a panda wings, and giving a donkey a turtle shell. In every example, our method successfully translates the 2D edit into geometry in a realistic manner. The edits in Figure 4 show a critical benefit of our method. Since the final mesh is constructed entirely from the output of a neural network, there are no geometric restrictions on the type of edit we can do. The last two rows demonstrate the ability of our network to change the genus of a shape by adding a handle or a hole through the middle, which would be impossible for geometry optimization-based methods.
@@
Figure 6: Mesh Editing Comparisons with Concurrent work: We compare our approach to concurrent work enabling localized 3D editing. While localized approaches better preserve the original structure of the shape, other methods are not able to produce edits as realistic as ours.
@@ 4.4 Ablations Studies & Discussion
Geometric Supervision. We investigate the effect of using geometric losses during training. Figure 7 compares three LRMs: a model trained by the pipeline described in Sec. 3, a model trained with no geometric supervision, and a model trained by replacing the normal map loss with a depth map loss. Using no geometric supervision results in poor surface quality, highlighted in the red boxes. Since the main training objective is multi-view image reconstruction, the model hallucinates correct geometry using colors, without producing an accurate surface. Supervising the predicted depth somewhat mitigates this issue, but the effect is weak and the surfaces are still incomplete. Normal map supervision gives high quality surfaces as shown in the green boxes.
Random Masking. We validate our choice in masking strategy by ablating our method using a uniformly random MAE-style mask across all views. This produces a clear train-test gap as during inference, we are always interested in editing contiguous 3D regions. This gap manifests in blurry and incorrect edits. We refer to Section B and Figure 2 of the supplementary material for details.
Table 2: Runtime Comparison: Our method is significantly faster than optimization methods as it is feed-forward and also faster than LRM-based approaches that must run multi-view diffusion.
Optimization-based LRM-based
TextDeformer MagicClay InstantMesh Instant3dit PrEditor3D Ours
Runtime $\downarrow$ $\scriptstyle\mathtt{\sim}$20mins $\scriptstyle\mathtt{\sim}$1hour 30sec $\scriptstyle\mathtt{\sim}$6sec 80sec <  3sec
Runtime: In Table 2 we provide performance comparisons between our approach and several top-performing recent works. Our method is not only much faster than optimization-based approaches [6, 27] as it requires only one forward pass, but it also outperforms LRM approaches [90, 5] that make use of a multi-view generation model. PrEditor3D [24] requires the forward pass of several large pre-trained models [69, 51] resulting in a longer runtime.
Limitations & Future Work: Our method is constrained by the expressiveness of editing in the canonical view. While text-to-image models can create a wide range of results, capturing a specific idea may require significant iteration. Our method is upper-bounded by the uniformity of the Marching Cubes triangulation, and the LRM reconstruction quality which makes performing edits that require extremely intricate details challenging. Blurry artifacts may arise when trying to reconstruct fine details (e.g. face) but we did not see such issues with shapes like chairs. MagicClay [6], manually freezes the un-edited geometry part but we designed a solution without such interventions. Future work could focus on improving localization by developing techniques to merge the existing triangulation with the edited output.
@@ 5 Conclusion
In this paper we introduced a new method to perform 3D shape editing. Our work builds upon the recent progress of LRMs by introducing a novel multi-view input masking strategy during training. Our LRM is trained to “inpaint” the masked region using a clean conditional viewpoint to reconstruct the missing information. During inference, a user may pass a single edited image as the conditional input, prompting our model to edit the existing shape in just one forward pass. We believe our method is an significant advancement in shape editing, allowing users to create accurate and controllable edits without 3D modeling expertise.
@@ Introduction
Masking Ablation
In order to justify our masking strategy, we train our model masking patches uniformly randomly instead of using 3D occlusions. Figure 9 compares meshes extracted from a model trained using our 3D masking versus masking $25\%$ of patches uniformly at random. We observe that uniformly random patch masking can still generate “roughly correct” shapes, especially adding a moustache to the face in the last row. This is because we add camera pose embeddings after masking, so the model can differentiate masked and non-masked tokens, regardless of their distribution within the image. Furthermore, the reconstruction outside of the masked region is still accurate. However, there still exists a train-test gap between random patches and contiguous patches created by selecting an editing region, which causes significant artifacts in the other three examples. In the first and second rows, we observe a blurring artifact, where the model cannot generate sharp features in the horns on the bird and between the slats of the chair. In the third row, using random patches causes the shape of the turtle shell to be malformed. In comparison, using our masking method produces accurate and sharp geometry in all examples.
Comparison to Tailor3D
@@
Mesh Editing Metrics
Large scale quantitative editing comparisons are difficult as there is no standard benchmark. In the Table above we compare against MagicClay [6] and Instant3Dit [5], generating edits using a text prompt. We then use the same text input for generating our conditional edited views and use CLIP similarity averaged across multiple views to measure the faithfulness of each method to the prompt.
Mesh Editing Comparisons
ViT-L-14 ViT-BigG-14
Ours Instant3Dit MagicClay Ours Instant3Dit MagicClay
CLIP Cos.Sim. $\uparrow$ 0.323 0.303 0.285 0.337 0.309 0.286
Table 3: Mesh Editing Comparisons: We provide CLIP cosine similarity metrics of our proposed MaskedLRM approach against two recent 3d editing techniques.
Additional Implementation Details
@@
We trained our models using 64 H100 GPUs with 80GB of RAM each. We use an AdamW optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$ and a weight decay of $0.01$. During stage 1 of training, we train for $30$ epochs. For each batch consisting of $12$ shapes, we randomly sample the number of input views for the batch uniformly at random between $6$ and $8$, not including the $1$ view for the conditional view. We use another $4$ views for supervision. Over the first 1500 iterations, we linearly warm up to a peak learning rate of 4e44e-4 and then use cosine learning rate decay. During stage 2, to account for increased rendering costs, we reduce the batch size to $8$ shapes. We train for $20$ epochs, with a peak learning rate of 5e65e-6.
Additional Qualitative Examples